xdx+ydy=a2(xdy-ydx)/x2+y2

1
Expert's answer
2021-05-07T12:01:27-0400

"xdx+ydy=\dfrac{a^2(xdy-ydx)}{x^2+y^2)}\\[9pt]

d(x^2+y^2)=\dfrac{a^2\frac{(xdy-ydx)}{x^2}}{1+\frac{y^2}{x^2}}\\[9pt]

d(x^2+y^2)=\dfrac{a^2d(\frac{y}{x})}{1+(\frac{y}{x})^2}\\[9pt]

\text{ Integrating Both the sides}\\[9pt]

\int d(x^2+y^2)=\int \dfrac{a^2d(\frac{y}{x})}{1+(\frac{y}{x})^2}\\[9pt]

\Rightarrow x^2+y^2=a^2arctan(\dfrac{y}{x})+C\\[9pt]

\Rightarrow x^2+y^2-a^2arctan(\dfrac{y}{x})=C"